An electric field of intensity 2.50 kN/C is applied along the x-axis. Calculate the electric flux through a rectangular plane 0.350 m wide and 0.700 m long if the following conditions are true (a) The plane is parallel to the yz-plane N m2/c (b) The plane is parallel to the xy-plane. N m2/c (c) The plane contains the y-axis, and its normal makes an angle of 32.5° with the x-axis. N m2/c
Answer #1

Solution)

We know, area of plane = 0.350 * 0.700 = 0.245 m2

(a) flux = E A cosφ = 2500 * 0.245 * cos0 = 612.5 Nm^2/C

======

(b) flux = 2500 * 0.350 *cos90 = 0

=======

(c) flux = 2500 * 0.350 *cos 32.5= 737.96

========